OpenStudy (anonymous):

Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used. All work must be shown. 1 over the cube root of the quantity of x to the negative sixth power

OpenStudy (anonymous):



OpenStudy (anonymous):

@DebbieG Please explain if u can

OpenStudy (debbieg):

Well first off, a negative power means to take the reciprocal. So \[\Large \dfrac{ 1 }{ \sqrt[n]{x^{-m}} }=\dfrac{ 1 }{ \sqrt[n]{x}^{-m} }= \sqrt[n]{x}^{m}\] And the root becomes the den'r in the rational exponent: \[\Large \sqrt[n]{x}^{m}=x^{\frac{m}{n}}\]

OpenStudy (anonymous):

Okay so u have to tur it over right

OpenStudy (anonymous):

 @DebbieG